# Optimizing Fine-Grained Parallelism Through Dynamic Load Balancing on Multi-Socket Many-Core Systems

### v1

@@ IV-C NUMA-aware Redirect Push (NA-RP)
In our first DLB strategy, NUMA-aware Redirect Push (NA-RP), the runtime redistribute tasks dynamically during work-sharing: see Alg. 3. The thief first repeats the logic in Alg. 1 $N_{\text{victim}}$ times—in effect, asking $N_{\text{victim}}$ victims to redirect tasks to it. Each victim, upon successfully handling a request using Alg. 2, calls doLoadBalancing(): see Alg. 3. doLoadBalancing() then changes victim state to be ready to redirect $N_{\text{steal}}$ new tasks to the thief: see doRedirectPush(). If the thief’s task queue is full or the victim handles all redirect push requests, the victim increments the round number and is ready to accept new request.
@@
The NA-RP strategy extends XQueue’s lock-less enqueuing mechanisms by only altering the target task queue to which new tasks are pushed. It prioritizes helping under-loaded workers to mitigate work imbalance. The additional overhead lies in the exchange of messages across cores.
@@ IV-D NUMA-aware Work Stealing (NA-WS)
We first explore a queue-based strategy. Recall that each XQueue thread possesses $N_{\text{worker}}\times S_{\text{queue}}$ task queues where $N_{\text{worker}}$ is the number of workers and $S_{\text{queue}}$ is the size of each individual queue. We allocate $N_{\text{worker}}$ request cells and round cells that map to each queue. The routine starts with the thief picking victims with $p_{\text{local}}$ of picking NUMA-local workers. It then randomly picks a queue in each of these victims, sending requests using our lock-less communication protocol. Upon receiving requests, a potential victim scans a subset at a time to find one to further process. This strategy guarantees that there is only one producer and consumer for each request cell, thus avoiding requests being overwritten, which could result in retry delays for the thieves. Through experiments, we observed that this method does not mitigate work imbalance and it introduces additional overhead. With millions of steal requests sent, only few tasks are stolen. We observed that 62% of requests are handled by the victims. However, less than 1% of all requests found by victims are valid, and around 0.01% of these valid requests result in successful steals. The runtime is rarely able to steal because there is a mismatch between the number of requests sent and the number of requests handled.
We further improve this strategy by letting a thief steal a batch of tasks from a victim, rather than requesting from a victim’s individual task queues: thus reducing communication from queue-to-queue to worker-to-worker. The thief uses the logic in Alg. 1. A victim, upon successfully handling a request, dequeues up to $N_{\text{steal}}$ tasks from its queue and en-queues them to the thief’s target task queue. This round of stealing completes when 1) no task is found from the victim; 2) the thief’s target task queue is full; or 3) $N_{\text{steal}}$ tasks have been migrated. We present the pseudo-code in Alg. 4.
@@
Our lock-less work stealing approach inherits both lock-less enqueuing and dequeuing mechanisms from XQueue by migrating already-queued tasks from one worker’s queue to another instead of co-locating the newly spawned tasks. Theoretically, this strategy incurs slightly more overhead compared to NA-RP because of the additional dequeuing operations. However, it could result in far fewer tasks stolen than tasks redirected in the NA-RP approach because stealing $N_{\text{steal}}$ tasks is the upper bound for each round of stealing, while redirecting $N_{\text{steal}}$ is the upper and the lower bound for each round of redirection. Unlike NA-RP that tends to push the tasks away from where they are created to remote workers, our NA-WS approach tends to bring the tasks back to where they are created, thus exploiting benefits of data proximity.
@@ IV-E DLB Configuration
We implement several configurable parameters to control how the DLB strategies operate. These parameters are the number of victims ($N_{\text{victim}}$) to whom a worker sends requests each time it becomes a thief; the max number of tasks to be stolen/redirected ($N_{\text{steal}}$) for each request; the timeout interval ($T_{\text{interval}}$) between two requests (addressing the situation that an under-loaded worker could remain idle for an extended period of time if requests sent while idle are not answered); and NUMA-local probability ($P_{\text{local}}$) controlling the probability that a thief steals from NUMA-local nodes.
@@ V Software Profiling Tools
To better analyze the performance characteristics of OpenMP tasking we developed new per-thread performance profiling tools. These tools enable us to track the time spent on tasking runtime operations. We categorize the operations that are most important for performance analysis into several event types. We identify the start and end of each event and insert marker functions perf_record with corresponding settings to record the event to profiler memory. We use rdtscp (ReaD Time-Stamp Counter and Processor) id value as the timestamp for each event. rdtscp is a light-weight processor intrinsic instruction that reads the current value of the processor’s timestamp counter. The processor increments the timestamp counter monotonically every clock cycle and resets it to 0 when the processor is reset. Note that while rdtscp is not a serializing instruction, it ensures that all prior instructions have executed and all previous loads are globally visible [13]. At the conclusion of the program, the xomp_perflog_dump API can be used to save the logs to the file system in a path defined by environment variables.
@@
Although enabling performance logging introduces overheads for fine-grained tasks, e.g., due to increased memory accesses when storing event information and the unavoidable hardware overhead of rdtscp, it still captures overall runtime behaviors, as each logging operation overhead is almost constant. Below, we describe our event classes. Each event class includes a pair of start and end timestamps.
@@
The cycles spent by a thread for different purposes are tracked as follows: By thread on the system: THREAD; by a task: TASK; creating tasks (crucial because fine-grained tasks can spend a large portion of their lifecycle on task creation): GOMP_TASK; waiting: TASKWAIT. BARRIER denotes the number of cycles spent on a barrier. If a thread is unoccupied, i.e., no task is scheduled in the current thread and it is busy checking its task queue, we record the cycles with STALL.
@@
We also instrument per-thread statistical counters. These counters incur little overhead as they are thread-local and can use lower-level caches. We implement a set of general performance counters as well as specialized counters for our DLB strategies. To track the locality of a task, we assign the thread ID to the task upon creation.
@@
To manage data locality, XGOMPTB tracks the number of tasks executed: i) by the thread that created it (NTASKS_SELF), ii) by the NUMA node that created it (NTASKS_LOCAL) iii) by another node (NTASKS_REMOTE). XGOMPTB also tracks the number of tasks that are i) statically pushed (NTASKS_STATIC_PUSH and ii) are not pushed due to the target queue being full and are executed immediately (NTASKS_IMM_EXEC).
@@
In addition, for the two DLB strategies we track the number of steal requests (NREQ_SENT), handled requests (NREQ_HANDLED), and stolen tasks (NTASKS_STOLEN). Handled requests are further categorized into requests that i) result in a task being stolen (NREQ_HAS_STEAL), or ii) fail due to an attempt to steal tasks from an empty queue (NREQ_SRC_EMPTY) or to a full queue (NREQ_TARGET_FULL). Stolen tasks are further categorized into tasks stolen by i) NUMA-local thieves or ii) NUMA-remote thieves.
@@ VI Evaluation
We evaluate our XGOMP and XGOMPTB implementations and NA-RP and NA-WS strategies using nine applications from the Barcelona OpenMP Task Suite (BOTS) [8]: Fibonacci (Fib), NQueens, Fast Fourier Transform (FFT), Floor Plan (FP), Health, Unbalance Tree Search (UTS), Strassen Matrix Multiplication (STRAS), and Protein Alignment (Align). We first compare XGOMP with the original GNU OpenMP (GOMP). We also include two LLVM variants—LLVM OpenMP (LOMP) and LLVM OpenMP using XQueue (XLOMP)—in our results for comparison. We conduct all experiments on an Intel Skylake-192 machine, with 192 cores (384 hardware threads) and eight NUMA zones. We compile the benchmarks with GNU GCC version 12.2.1 and LLVM version 14.0.0, both with -fopenmp flag and O3 optimization level. We use up to 192 threads and bind OpenMP threads to cores with close thread affinity. We use the following input arguments for our applications. Fib: 42, NQueens: 15, FFT: 536M, FP: 20, Health: large, UTS: small, STRAS: 8192 (Y=16), Sort: 1B, and Align: 1000. For our DLB experiments (Section VI-B), in which we run large parameter sweeps over many configuration parameters, we scale down Fib: 40, FFT: 268M and Sort: 268M, FP: 15, UTS: tiny, and STRAS: 2048 (Y=16) to reduce experiment times. We have validated experimentally that the larger and smaller experiments yield similar results.
@@ VI-A GNU OpenMP with XQueue and Distributed Tree Barrier
To understand these results, we measure the time spent on various runtime profiling events (see Section V). We observe that, applications for which LOMP and XLOMP outperform GOMP and XGOMP—Fib, NQueens, FP, Health, and UTS—spend much of thread time on task creation (mostly task allocation). While applications that XGOMP and XGOMPTB outperform XLOMP—FFT, STRAS, Sort, and Align—spend most of their time on task execution. These results are due to the way that memory is allocated. All OpenMP implementations use malloc to allocate memory for tasks. malloc is thread safe and thus uses synchronization. The default allocation approach used by Linux prioritizes allocating memory in the NUMA node closest to the CPU executing the task. GOMP invokes malloc each time it creates a new task, whereas LOMP instead allocates memory using a fast, multi-level allocator, which pre-allocates chunks of buffer space and manages this space itself. When a thread creates a new task, LOMP/XLOMP will ii) use a local buffer if available; ii) synchronously but locality-agnostically acquire a buffer from another thread if available; or iii) call malloc. For fine-grained tasks, large portions of time are spent on task allocation/de-allocation. For GOMP, threads contend for malloc operations, which leads to sequential mallocs. LOMP can parallelize memory allocation, because method i) is frequently used. For large tasks, LOMP/XLOMP performs a similar number of malloc requests to GOMP as the multi-level allocator is less beneficial. That is, methods i) and ii) are less likely to be successful (as longer tasks consume the buffer for longer periods). Further, LOMP “steals” buffer space using method ii), which results in more tasks executed remotely. LOMP has slightly higher task execution overhead because it handles a richer set of cases.
@@ VI-B XGOMPTB with Dynamic Load Balancing
We study our two lock-less DLBs with the same BOTS benchmarks using the Skylake-192 machine. We conduct experiments with 192 threads on eight NUMA zones, controlling runtime behavior with the configuration parameters described in Section IV-E. For each application, we conduct a parameter sweep to identify the optimal settings for each value of $N_{\text{victim}}$, $N_{\text{steal}}$, $T_{\text{interval}}$ and $P_{\text{local}}$. We run each DLB configuration ten times, and compare its average execution time with XGOMPTB.
@@ VII Application in Blockchain Consensus: Proof-of-Space (PoSp)
The state-of-the-art production PoSp blockchains (Chia  [20]) use a minimum of $2^{K}$ (where $K=32$) cryptographic puzzles to be stored in a single file [17, 21]. The PoSp evaluated is made up of $2^{K}$ cryptographic puzzles, where each puzzle is composed of a 28 Bytes BLAKE3 hash and a 4 Bytes nonce value. We conducted a parameter sweep (see Fig. 8) comparing XGOMPTB and GOMP on our 192-core system with increasing batch size. Using a batch size of 1, XGOMPTB achieves 7.8 megahashes per second (MH/s), compared to 0.04 MH/s for native GOMP, a 195$\times$ improvement. A batch size of 1 yields relatively small tasks, stressing the runtime’s ability to process a high throughput of tasks per second. XGOMPTB achieves 7.8 M tasks per second while GOMP only achieves 40 K tasks per second. The significant throughput reduction for GOMP comes from shared locks in the OpenMP runtime. The best performance is achieved for a batch size of 1024 for XGOMPTB with 217 MH/s, compared to 164 MH/s for GOMP with an 8192 batch size, a 32% speedup. Batch sizes that are too large can yield lower performance stemming from load imbalance that leads to underutilized resources.
@@ VIII Performance Tuning
To further study relationships between the performance and parameters, we categorize the applications into five sizes based on per-task rdtscp cycles $S_{\text{task}}$. We then define steal size $S_{\text{steal}}$ in Equation 1, where $N_{\text{victim}}$ is the number of victims per request, $N_{\text{steal}}$ is the number of steals per victim, and $T_{\text{interval}}$ is a timeout counter for the worker before it retries next round of requests. Since applications lack sensitivity to $T_{\text{interval}}$, we reduce its influence by applying a $\log$ function.
@@
Fig. 10 shows that NA-WS is less sensitive to configuration and leads to performance degradation only for applications with small task size ($<$103) and large steal size. As task size increases, application performance generally increases, with the best improvement achieved for the largest task and steal sizes. Generally, applications with larger task sizes benefit more from larger steal sizes; applications with small task size should use small steal sizes.
@@
Locality choice is also an important contributor to performance. For NA-RP, fully local redirection yields the best performance across all benchmarks (see Table I) because it optimizes memory access latency by pushing tasks to adjacent workers. Because NA-RP pushes more tasks away, it is the best choice to send tasks to the local node which can also utilize multi-level cache. For NA-WS, extremely fine-grained tasks perform best with fully NUMA-local steal, to optimize the overall memory access latency. For larger tasks, smaller $P_{\text{local}}$ performs better because work stealing can return pushed tasks to where they are created. Therefore, tasks may utilize high-level cache which optimizes overall latency.
In summary, different combinations of strategies and parameters fit different application characteristics. Applications with extremely fine-grained tasks benefit from smaller steal sizes on the NA-WS approach; full NUMA-local probability should be used. Applications with larger tasks benefit from larger steal sizes. NA-RP with maximum steal size and fully NUMA-local is particularly suitable for tasks of $>$104 cycles. We summarize in Table IV guidelines for selecting parameters for the best performance and show in Fig. 11 results obtained for the five applications when using these guidelines. We use as input arguments, Fib: 42, NQueens: 16, FFT: 536M, FP: 20, Health: xlarge, UTS: small, STRAS: 4096 (Y=16), Sort: 1B, and Align: 2000.

### v2

@@ IV-C NUMA-aware Redirect Push (NA-RP)
In our first DLB strategy, NUMA-aware Redirect Push (NA-RP), the runtime redistributes tasks dynamically during work-sharing: see Alg. 3. The thief first repeats the logic in Alg. 1 $N_{\text{victim}}$ times—in effect, asking $N_{\text{victim}}$ victims to redirect tasks to it. Each victim, upon successfully handling a request using Alg. 2, calls doLoadBalancing(): see Alg. 3. doLoadBalancing() then changes victim state to be ready to redirect $N_{\text{steal}}$ new tasks to the thief: see doRedirectPush(). If the thief’s task queue is full or the victim handles all redirect push requests, the victim increments the round number and is ready to accept new request.
@@
The NA-RP strategy extends XQueue’s lock-less enqueuing mechanisms by only altering the target task queue to which new tasks are pushed. It prioritizes helping underloaded workers to mitigate work imbalance. The additional overhead lies in the exchange of messages across cores.
@@ IV-D NUMA-aware Work Stealing (NA-WS)
We first explore a queue-based strategy. Recall that each XQueue thread possesses $N_{\text{worker}}\times S_{\text{queue}}$ task queues where $N_{\text{worker}}$ is the number of workers and $S_{\text{queue}}$ is the size of each individual queue. We allocate $N_{\text{worker}}$ request cells and round cells that map to each queue. The routine starts with the thief picking victims with $p_{\text{local}}$ of picking NUMA-local workers. It then randomly picks a queue in each of these victims, sending requests using our lock-less communication protocol. Upon receiving requests, a potential victim scans a subset at a time to find one to further process. This strategy guarantees that there is only one producer and consumer for each request cell, thus avoiding requests being overwritten, which could result in retry delays for the thieves. Through experiments, we observed that this method does not mitigate work imbalance and it introduces additional overhead. With millions of steal requests sent, only few tasks are stolen. We observed that 62% of requests are handled by victims. However, less than 1% of all requests found by victims are valid, and around 0.01% of these valid requests result in successful steals. The runtime is rarely able to steal because there is a mismatch between the number of requests sent and the number of requests handled.
We further improve this strategy by letting a thief steal a batch of tasks from a victim, rather than requesting from a victim’s individual task queues: thus reducing communication from queue-to-queue to worker-to-worker. The thief uses the logic in Alg. 1. A victim, upon successfully handling a request, dequeues up to $N_{\text{steal}}$ tasks from its queue and enqueues them to the thief’s target task queue. This round of stealing completes when 1) no task is found from the victim; 2) the thief’s target task queue is full; or 3) $N_{\text{steal}}$ tasks have been migrated. We present the pseudo-code in Alg. 4.
@@
Our lock-less work stealing approach inherits both lock-less enqueuing and dequeuing mechanisms from XQueue by migrating already-queued tasks from one worker’s queue to another instead of co-locating the newly spawned tasks. Theoretically, this strategy incurs slightly more overhead compared to NA-RP because of the additional dequeuing operations. However, it could result in far fewer tasks stolen than tasks redirected in the NA-RP approach because stealing $N_{\text{steal}}$ tasks is the upper bound for each round of stealing, while redirecting $N_{\text{steal}}$ is the upper and lower bound for each round of redirection. Unlike NA-RP, which tends to push tasks away from where they are created to remote workers, NA-WS tends to bring the tasks back to where they are created, thus exploiting benefits of data proximity.
@@ IV-E DLB Configuration
We implement several configurable parameters to control how the DLB strategies operate. These parameters are the number of victims ($N_{\text{victim}}$) to whom a worker sends requests each time it becomes a thief; the max number of tasks to be stolen/redirected ($N_{\text{steal}}$) for each request; the timeout interval ($T_{\text{interval}}$) between two requests (addressing the situation that an underloaded worker could remain idle for an extended period of time if requests sent while idle are not answered); and NUMA-local probability ($P_{\text{local}}$) controlling the probability that a thief steals from NUMA-local nodes.
@@ V Software Profiling Tools
To better analyze the performance characteristics of OpenMP tasking we developed new per-thread performance profiling tools. These tools enable us to track the time spent on tasking runtime operations. We categorize the operations that are most important for performance analysis into several event types. We identify the start and end of each event and insert marker functions perf_record with corresponding settings to record the event to profiler memory. We use rdtscp (ReaD Time-Stamp Counter and Processor) id value as the timestamp for each event. rdtscp is a light-weight processor intrinsic instruction that reads the current value of the processor’s timestamp counter. The processor increments the timestamp counter monotonically every clock cycle and resets it to 0 when the processor is reset. Note that while rdtscp is not a serializing instruction, it ensures that all prior instructions have executed and all previous loads are globally visible [13]. At the end of the program, the xomp_perflog_dump API can be used to save the logs to the file system in a path defined by environment variables.
@@
Although enabling performance logging introduces overheads for fine-grained tasks, for example, due to increased memory accesses when storing event information and the unavoidable hardware overhead of rdtscp, it still captures overall runtime behaviors, as each logging operation overhead is almost constant. Below, we describe our event classes. Each event class includes a pair of start and end timestamps.
@@
The cycles spent by a thread for different purposes are tracked as follows: By a task: TASK; creating tasks (crucial because fine-grained tasks can spend a large portion of their lifecycle on task creation): GOMP_TASK; and waiting: TASKWAIT. BARRIER denotes the number of cycles spent on a barrier. If a thread is unoccupied, i.e., no task is scheduled in the current thread and it is busy checking its task queue, we record the cycles as STALL.
@@
We also instrument per-thread statistical counters. These counters incur little overhead as they are thread-local and can use lower-level caches. We implement a set of general performance counters and specialized counters for our DLB strategies. To track the locality of a task, we assign the thread ID to the task upon creation.
@@
To manage data locality, XGOMPTB tracks the number of tasks executed: i) by the thread that created it (NTASKS_SELF); ii) by the NUMA node that created it (NTASKS_LOCAL); and iii) by another node (NTASKS_REMOTE). XGOMPTB also tracks the number of tasks that are i) statically pushed (NTASKS_STATIC_PUSH and ii) are not pushed due to the target queue being full and are instead executed immediately (NTASKS_IMM_EXEC).
@@
In addition, for the two DLB strategies we track the number of steal requests (NREQ_SENT), handled requests (NREQ_HANDLED), and stolen tasks (NTASKS_STOLEN). Handled requests are further categorized into requests that i) result in a task being stolen (NREQ_HAS_STEAL); or ii) fail due to an attempt to steal tasks from an empty queue (NREQ_SRC_EMPTY) or to a full queue (NREQ_TARGET_FULL). Stolen tasks are further categorized into tasks stolen by i) NUMA-local thieves or ii) NUMA-remote thieves.
@@ VI Evaluation
We evaluate our XGOMP and XGOMPTB implementations and NA-RP and NA-WS strategies using nine applications from the Barcelona OpenMP Task Suite (BOTS) [8]: Fibonacci (Fib), NQueens, Fast Fourier Transform (FFT), Floor Plan (FP), Health, Unbalance Tree Search (UTS), Strassen Matrix Multiplication (STRAS), and Protein Alignment (Align). We first compare XGOMP with the original GNU OpenMP (GOMP). We also include two LLVM variants—LLVM OpenMP (LOMP) and LLVM OpenMP using XQueue (XLOMP)—in our results for comparison. We conduct all experiments on an Intel Skylake-192 machine, with 192 cores (384 hardware threads) and eight NUMA zones. We compile the benchmarks with GNU GCC version 12.2.1 and LLVM version 14.0.0, both with the -fopenmp flag and O3 optimization level. We use up to 192 threads and bind OpenMP threads to cores with close thread affinity. We use the following input arguments for our applications. Fib: 42, NQueens: 15, FFT: 536M, FP: 20, Health: large, UTS: small, STRAS: 8192 (Y=16), Sort: 1B, and Align: 1000. For our DLB experiments (Section VI-B), in which we run large parameter sweeps over many configuration parameters, we scale down Fib: 40, FFT: 268M and Sort: 268M, FP: 15, UTS: tiny, and STRAS: 2048 (Y=16) to reduce experiment times. We have validated experimentally that the larger and smaller experiments yield similar results.
@@ VI-A GNU OpenMP with XQueue and Distributed Tree Barrier
To understand these results, we measure the time spent on various runtime profiling events (see Section V). We observe that applications for which LOMP and XLOMP outperform GOMP and XGOMP—Fib, NQueens, FP, Health, and UTS—spend much of thread time on task creation (mostly task allocation). While applications for which XGOMP and XGOMPTB outperform XLOMP—FFT, STRAS, Sort, and Align—spend most of their time on task execution. These results are due to the way that memory is allocated. All OpenMP implementations use malloc to allocate memory for tasks. malloc is thread safe and thus uses synchronization. The default allocation approach used by Linux prioritizes allocating memory in the NUMA node closest to the CPU executing the task. GOMP invokes malloc each time it creates a new task, whereas LOMP instead allocates memory using a fast multi-level allocator, which pre-allocates chunks of buffer space and manages this space itself. When a thread creates a new task, LOMP/XLOMP will i) use a local buffer if available; ii) synchronously but locality-agnostically acquire a buffer from another thread if available; or iii) call malloc. For fine-grained tasks, large portions of time are spent on task allocation/de-allocation. For GOMP, threads contend for malloc operations, which leads to sequential mallocs. LOMP can parallelize memory allocation, because method i) is frequently used. For large tasks, LOMP/XLOMP performs a similar number of malloc requests to GOMP as the multi-level allocator is less beneficial. That is, methods i) and ii) are less likely to be successful (as longer tasks consume the buffer for longer periods). Further, LOMP “steals” buffer space using method ii), which results in more tasks executed remotely. LOMP has slightly higher task execution overhead because it handles a richer set of cases.
@@ VI-B XGOMPTB with Dynamic Load Balancing
We study our two lock-less DLBs with the same BOTS benchmarks using the Skylake-192 machine. We conduct experiments with 192 threads on eight NUMA zones, controlling runtime behavior with the configuration parameters described in Section IV-E. For each application, we conduct a parameter sweep to identify the optimal settings for each value of $N_{\text{victim}}$, $N_{\text{steal}}$, $T_{\text{interval}}$ and $P_{\text{local}}$. We run each DLB configuration ten times and compare its average execution time with XGOMPTB.
@@ VII Application in Blockchain Consensus: Proof-of-Space (PoSp)
The state-of-the-art production PoSp blockchains (e.g., Chia [20]) use a minimum of $2^{K}$ (where $K=32$) cryptographic puzzles that are stored in a single file [17, 21]. The PoSp we evaluated is made up of $2^{K}$ cryptographic puzzles, where each puzzle is composed of a 28 Bytes BLAKE3 hash and a 4 Bytes nonce value. We conducted a parameter sweep (see Fig. 8) comparing XGOMPTB and GOMP on our 192-core system with increasing batch size. Using a batch size of 1, XGOMPTB achieves 7.8 megahashes per second (MH/s), compared to 0.04 MH/s for native GOMP, a 195$\times$ improvement. A batch size of 1 yields relatively small tasks, stressing the runtime’s ability to process a high throughput of tasks per second. XGOMPTB achieves 7.8 M tasks per second while GOMP only achieves 40 K tasks per second. The significant throughput reduction for GOMP comes from shared locks in the OpenMP runtime. The best performance is achieved for a batch size of 1024 for XGOMPTB with 217 MH/s, compared to 164 MH/s for GOMP with an 8192 batch size, a 32% speedup. Batch sizes that are too large can yield lower performance due to load imbalance that leads to underutilized resources.
@@ VIII Performance Tuning
To further study the relationships between the performance and parameters, we categorize applications into five sizes based on per-task rdtscp cycles $S_{\text{task}}$. We then define steal size $S_{\text{steal}}$ in Equation 1, where $N_{\text{victim}}$ is the number of victims per request, $N_{\text{steal}}$ is the number of steals per victim, and $T_{\text{interval}}$ is a timeout counter for the worker before it retries the next round of requests. Since applications lack sensitivity to $T_{\text{interval}}$, we reduce its influence by applying a $\log$ function.
@@
Fig. 10 shows that NA-WS is less sensitive to configuration and leads to performance degradation only for applications with small tasks ($<$103) and large steal size. As task size increases, application performance generally increases, with the best improvement achieved for the largest task and steal sizes. Generally, applications with larger tasks benefit more from larger steal sizes; applications with small tasks should use small steal sizes.
@@
Locality is also an important contributor to performance. For NA-RP, fully local redirection yields the best performance across all benchmarks (see Table I) because it optimizes memory access latency by pushing tasks to adjacent workers. Because NA-RP pushes more tasks away, it is the best choice to send tasks to the local node, which can also utilize multi-level cache. For NA-WS, extremely fine-grained tasks perform best with fully NUMA-local steal, to optimize the overall memory access latency. For larger tasks, smaller $P_{\text{local}}$ performs better because work stealing can return pushed tasks to where they are created, allowing them to utilize high-level cache which optimizes overall latency.
In summary, different combinations of strategies and parameters suit different application characteristics. Applications with extremely fine-grained tasks benefit from smaller steal sizes on the NA-WS approach; full NUMA-local probability should be used. Applications with larger tasks benefit from larger steal sizes. NA-RP with maximum steal size and fully NUMA-local is particularly suitable for tasks of $>$104 cycles. We summarize in Table IV guidelines for selecting parameters for the best performance and show in Fig. 11 results obtained for the five applications when using these guidelines. We use as input arguments, Fib: 42, NQueens: 16, FFT: 536M, FP: 20, Health: xlarge, UTS: small, STRAS: 4096 (Y=16), Sort: 1B, and Align: 2000.